OpenStudy (anonymous):

Multiply and simplify 1 3/12*2 2/3

OpenStudy (jdoe0001):

\(\bf 1\cfrac{3}{12}\cdot 2\cfrac{2}{3}\implies \cfrac{1\cdot 12+3}{12}\cdot \cfrac{2\cdot 3+2}{3}\)

OpenStudy (anonymous):

15/12*8/3

OpenStudy (jdoe0001):

\(\bf 1\cfrac{3}{12}\cdot 2\cfrac{2}{3}\implies \cfrac{1\cdot 12+3}{12}\cdot \cfrac{2\cdot 3+2}{3}\implies \cfrac{15}{12}\cdot \cfrac{8}{3}\implies \cfrac{15\cdot 8}{12\cdot 3}\)

OpenStudy (anonymous):

120/36

OpenStudy (jdoe0001):

yeap now \(\bf \cfrac{15\cdot 8}{12\cdot 3}\implies \cfrac{120}{36}\\ \quad \\ {\color{blue}{ 120\implies 12\cdot 10\qquad 36\implies 12\cdot 3}}\\ \quad \\ \cfrac{120}{36}\implies \cfrac{\cancel{12}\cdot 10}{\cancel{12}\cdot 3}\)

OpenStudy (anonymous):

now what

OpenStudy (jdoe0001):

well, that's the simplified version the numerator and denominator, don't have any others factors in common to cancel out

OpenStudy (jdoe0001):

3 is prime, only factor is 1 and self 10 factors are 5 and 2 and 1 and self..... so no dice on more common factors above and below

OpenStudy (anonymous):

8 5/12

OpenStudy (anonymous):

A.8 5/12 B.9(-7/12) c.9 7/12 d.8 7/12 e. 8(-7/12)